A hovercraft stays off the ground by blowing air out underneath it. It can accelerate (forwards or backwards) at 0,6 m/s2 by blowing air with a steering fan (backwards or forwards). If the hovercraft is moving forwards at 12 m/s and the steering fan is suddenly turned to accelerate the hovercraft backwards, calculate:

a) the velocity of the hovercraft after 2 s

b) the time it takes for the hovercraft to reach zero velocity 

c) the velocity of the hovercraft after 29 s

d) the displacement of the craft when it has zero velocity

e) the displacement of the hovercraft when it is moving backwards at 16m/s

f) Clearly show what the value of h(max),t(total) and v

State whether the hovercraft is going backwards or forwards after each question’s answer.

(Make forward direction positive and backwards direction negative)




1
Expert's answer
2021-08-17T16:54:02-0400


Part (a)

we have to find velocity after 2sec is:

"v=u+at\\

v=12-0.6*2=10.8m/s"

(here a=-0.6m/s^2 because the hovercraft suddenly start moving backward)

Part (b)

we have to find time taken by hovercraft when its velocity become zero:

using same formula:

"v=0,u=12m\/s,a=-0.6m\/s^2"

"so, t=20sec"


Part (c)

velocity after 29sec is:

"v=12-(0.6*29)=2.6m\/s"


Part (d)

displacement when velocity becomes zero:

"v^2=u^2+2as\\

s=\frac{(u^2)}{2a}=\frac{(12*12)}{(2*.6)}=120m"


Part (e)

displacement when it is moving backward at velocity 16m/s

"so s=\frac{(v^2-u^2)}{2a}\\

s=\frac{(16*16-12*12)}{(2*0.6)}=93.3m"


Part(f)

"v=u+at" (1st eqaution of motion)

"v=0,u=12m\/s and a=-0.6"

therefore,"t=20sec" (time of going upward)

hence, total time of flight is "2*20=40 sec"

"h_{max}=\frac{(u^2)}{2a}=\frac{(12*12)}{(2*.6)}=120m"

"v= \frac{s}{t}=\frac{120}{40}= 3 m/s"